The area of the square above is 114 cm2. What is the length of a side of the square?
Answer by Maths68(1474)

Area of a square = length of a side * length of a side side
a=s*s
a = s^2
144=s^2
Take square root both sides of the equation
sqrt(144)=sqrt(s^2)
sqrt(12*12)=sqrt(s*s)
12=s
s=12
Length of side = 12 cm